A 6.0 volt battery is connected to a 3.0 ohm resistor for 1 hour. A) Find the current through the resistor. B) Find the total charge that passes through the resistor.

A) Find the power output of the battery in 22.1.  B) Find the power used by the resistor in 22.1.
Answer #1

A)

Current, I = V/R = 6/3 = 2 A

B)

Charge, q = It = 2 x 3600 = 7200 C

C)

Power, P = 6 x 2 = 12 W

D)

Power, P = 6^2/3 = 12 W

Comment in case any doubt please rate my answer...